# LeanCat: A Benchmark Suite for Formal Category Theory in Lean (Part I: 1-Categories)

### v1

@@ Abstract
Large language models (LLMs) have made rapid progress in formal theorem proving, yet current benchmarks under-measure the kind of abstraction and library-mediated reasoning that organizes modern mathematics. In parallel with FATE’s emphasis on frontier algebra, we introduce LeanCat [^1] LeanCat (Part I: 1-Categories) is open sourced at https://github.com/sciencraft/LeanCat., a Lean benchmark for category-theoretic formalization—a unifying language for mathematical structure and a core layer of modern proof engineering—serving as a stress test of structural, interface-level reasoning. Part I: 1-Categories contains 100 fully formalized statement-level tasks, curated into topic families and three difficulty tiers via an LLM-assisted + human grading process. The best model solves $8.25\%$ of tasks at pass@1 ($32.50\%/4.17\%/0.00\%$ by Easy/Medium/High) and $12.00\%$ at pass@4 ($50.00\%/4.76\%/0.00\%$). We also evaluate LeanBridge which use LeanExplore to search Mathlib, and observe consistent gains over single-model baselines. LeanCat is intended as a compact, reusable checkpoint for tracking both AI and human progress toward reliable, research-level formalization in Lean.
@@ 1 Introduction
Recent advances in large language models (LLMs) and agentic training have revived the prospect of end-to-end formal theorem proving. On the formal side, systems such as OpenAI’s early Lean-based prover (Openai2022) and DeepMind’s AlphaProof (DeepMind2025) demonstrate that reinforcement learning with formal verification feedback can produce non-trivial Lean proofs. More recently, specialized provers such as Seed-Prover 1.5 (SeepProver1.5) show that large-scale agentic RL and test-time scaling can push formal success rates on established benchmarks substantially higher. These results suggest that formal proof generation is no longer limited to toy domains, and that tight tool feedback loops (retrieve–generate–verify) can be a decisive ingredient.
@@
Despite steady progress in neural theorem proving, current formal benchmarks still do not adequately exercise library-grounded, abstraction-heavy reasoning. Widely used datasets such as miniF2F (miniF2F) and FIMO FIMO largely draw from olympiad-style problems, while university-level suites like ProofNet proofNet and PutnamBench putnamBench focus on undergraduate competition or textbook material. These benchmarks are valuable, but they often reward short clever tricks or computation rather than sustained work inside rich abstract frameworks. By contrast, modern research mathematics is written at high generality, organized around reusable interfaces, and deeply intertwined with large libraries of definitions and lemmas—so success hinges less on a single key insight and more on navigating abstraction, managing definitions, and coherently composing library knowledge over long horizons.
@@
Category theory provides a natural stress-test for this capability: as the interface language of modern mathematics—categories, functors, natural transformations, adjunctions, limits/colimits, monads—formal proofs typically rely on diagrammatic and universal-property reasoning, i.e., constructing morphisms with the right naturality/uniqueness guarantees and showing commutativity across families of structures. Yet existing formal benchmarks rarely target this abstraction layer directly. To bridge this gap, we introduce LeanCat, a benchmark of 100 category-theory problems formalized in Lean 4 (mathlib), designed to test whether automated provers can operate inside a mature library and compose high-level abstractions rather than solve isolated puzzles. LeanCat complements algebra-focused benchmarks such as the FATE series Fate by shifting the frontier from abstract algebra to category theory. Our baseline evaluation reveals a stark abstraction gap: across five strong models, the best achieves only $8.25\%$ at pass@1 and $12\%$ at pass@4, with accuracy dropping sharply from Easy to High once library navigation and long-horizon abstraction management become necessary (Figure 1). We further observe a persistent gap between producing a plausible natural-language argument and generating compilable Lean code, highlighting a pronounced natural-to-formal bottleneck (Figure 2).
Figure 1: Formal proof accuracy (pass@4) across LeanCat model baselines.
Figure 2: Intermediate natural-language (NL) vs. formal-language (FL) (pass@4) accuracy on LeanCat, highlighting the natural-to-formal gap.
@@
To our knowledge, LeanCat is the first installment of a category-theory benchmark series. This paper focuses on on 1-category theory. We envision future extensions targeting richer structures such as monoidal categories, enriched and braided settings, and eventually 2-categories and higher-categorical interfaces. Beyond benchmarking, we view this direction as meaningful for both (i) human mathematics, by mapping what parts of abstract library-level reasoning remain hard to formalize and where libraries need strengthening, and (ii) AI, by forcing progress on abstraction-aware planning, retrieval of relevant lemmas, and robust compilation-driven refinement.
@@
  • LeanCat benchmark (1-Categories). We present LeanCat, a set of 100 category-theory problems formalized in Lean 4 (mathlib 4.19.0). The tasks span eight topic clusters (from basic categorical properties to monads), curated to cover reusable abstract interfaces rather than contest-style tricks.
  • Difficulty annotation pipeline. We propose a grading process combining model-based estimates and expert judgment. Each problem receives multiple 1–10 ratings (from advanced LLM attempts and from human formalizers), aggregated with heavier weight on human scores to assign Easy/Medium/High labels (20/42/38).
  • Baseline evaluation. We benchmark state-of-the-art provers on LeanCat under uniform conditions. Our evaluation (Section 3-4) includes ChatGPT-5.1 and ChatGPT-5.2 (ChatGPT5.1; ChatGPT5.2), Claude 4.5 (ClaudeOpus4.5), Google’s Gemini 3 pro (Gemini3pro), an advanced chain-of-thought reasoner DeepSeek -V3.2-Thinking and DeepSeek V3.2 Speciale (DeepseekV3.2) and ann agentic model Kimi K2 KimiK2. At pass@1, the best model solves 8.25% of LeanCat; at pass@4, the best reaches 12%. We provide breakdowns by difficulty and identify dominant failure modes (library gaps, abstraction mismatch, and stalled multi-step plans).
  • Search-augmented proving via LeanBridge. We evaluate a retrieve–analyze–generate–verify loop that integrates mathlib retrieval (via LeanExplore) and compiler feedback, illustrating how tool-augmented workflows can improve robustness on a subset of problems.
@@ 2.1 Benchmark Structure and Content
  • Abstract part: Problems are primarily drawn from standard, widely used textbooks in category theory, especially Category Theory in Context (Riehl2017) and Categories for the Working Mathematician (ML1998), together with a small number of problems adapted from unpublished lecture notes (Kong; Zheng).
  • Concrete part: Problems are curated mainly from Abstract and Concrete Categories (AHS1990), which provides a systematic source of exercises on concretizability, injectivity, and related themes.
  • Others: Beyond these core sources, we also include selected problems inspired by research papers and advanced community-driven references (Chen2021; ACMU2021).
Each LeanCat problem is self-contained at the statement level: the formal statement of the theorem is provided (often with informal description as in the Problem listing above), and all necessary definitions exist in Lean’s environment (either already in Mathlib (mathlib) or introduced as part of the problem setup). Where possible, we drew on known theorems from category theory literature, many tasks were crafted or adapted specifically to probe corner cases and interface interactions that an AI prover might struggle with. In several High-difficulty cases, relevant lemmas were not readily available, forcing human formalizers to devise intermediate results. This aspect of LeanCat, that it sometimes goes beyond the library, makes it a particularly stringent test for automated provers, which cannot rely solely on rote application of existing library facts.
LeanCat comprises 100 theorem statements in category theory, each fully formalized in Lean 4 (i.e. each problem is given as a Lean theorem declaration, with required definitions and context available). The problems are organized into eight topical clusters reflecting key areas of category theory:
Figure 3: Sunburst diagram showing the distribution of our LeanBridge problem sets by topic and difficulty. The inner ring groups problems into thematic sections (Basic, Adjunction, Reflective, Concrete, Limit, Cocompletion, Abelian, Monad), while the outer ring lists individual problems, labelled by difficulty: E = Easy, M = Medium, H = High.
@@ 2.2 Curation Workflow
LeanCat is a statement-level benchmark: each item consists of a single standalone theorem to prove, rather than a scaffolded sequence of intermediate lemmas that gradually builds toward a final goal. This design aims to evaluate general library-grounded proving capabilities—retrieval, definition management, and abstraction navigation—instead of rewarding problem-specific hint engineering.
@@
All benchmark files follow strict, uniform conventions: (i) each Lean file contains exactly one sorry after the final theorem; (ii) the natural-language problem description (in ) is included as a comment immediately preceding the formal statement; (iii) universe levels are explicitly fixed to avoid ambiguity and instability that are common in category-theory developments.
@@ 2.3 Difficulty Annotation Pipeline
  • LLM Difficulty Scoring: For each model, a correct proof contributes a proof score, and (only when that model does not already have a correct proof) a correct statement contributes a smaller statement score. The total score of a problem is the weighted sum of these contributions across models; the difficulty is then defined as Diff==max (0,10-PF score-ST score), so that problems solved by none of the models have difficulty 10, while problems for which all proof columns are green have difficulty 0.
  • Human Difficulty Scoring: In parallel, two human mathematicians (with Lean expertise and category theory background) independently rated each problem on the same 1-10 difficulty scale. They factored in things like the length of the proof, intricacy of arguments, and whether any non-obvious lemmas are required. The human scores tended to correlate with intuition: e.g. a trivial diagram chase might be 2/10, whereas a complex construction spanning several definitions could be 9/10.
  • Aggregation: We combined the scores giving 50% weight to human ratings and 50% to LLM ratings. Finally, we mapped the numeric scores to difficulty categories. We defined thresholds informed by the distribution of scores: roughly, Easy for scores $\leq$6, High for scores $\geq$8.5, and Medium for the rest. These cut-offs neatly split the set into 20 Easy, 42 Medium, and 38 High, see Table 4.
Figure 4: We use LeanBridge to benchmark LeanCat across several LLMs. The table records, for each problem and model on natural-language statements, whether the model produced a correct statement or proof. A green cell indicates that the model produced a correct Lean proof of the intended statement. A blue cell indicates that the model produced a correct statement but without a correct proof. An olive cell indicates that the model’s statement is not phrased in the same natural language as the original problem. A grey cell indicates that the output is not a complete Lean script (for example, it contains syntax or type errors).
This joint annotation procedure yields richer insight than a single-expert classification. It effectively uses large models as “second-opinion graders.” The resulting difficulty labels have already proved useful in analysis: for instance, all of the seven problems solved by the best model (Section 4) came from the Easy set, and problems with score $\geq$9 (the “hardest of High”) saw zero successes across all models - quantitative evidence that our difficulty rankings align with actual solvability.

### v2

@@ Abstract
While large language models (LLMs) have demonstrated impressive capabilities in formal theorem proving, current benchmarks fail to adequately measure library-grounded abstraction—the ability to reason with high-level interfaces and reusable structures central to modern mathematics and software engineering. We introduce LeanCat, a challenging benchmark comprising 100 fully formalized category-theory tasks in Lean. Unlike algebra or arithmetic, category theory serves as a rigorous stress test for structural, interface-level reasoning. Our evaluation reveals a severe abstraction gap: the best state-of-the-art model solves only $12.0\%$ of tasks at pass@4, with performance collapsing from $55.0\%$ on Easy tasks to $0.0\%$ on High-difficulty tasks, highlighting a failure in compositional generalization. To overcome this, we evaluate LeanBridge, a retrieval-augmented agent that employs a retrieve–generate–verify loop. LeanBridge achieves a peak success rate of $24.0\%$—doubling the performance of the best static baseline. These results empirically demonstrate that iterative refinement and dynamic library retrieval are not merely optimizations but strict necessities for neuro-symbolic reasoning in abstract domains. LeanCat offers a compact, reusable testbed for tracking progress toward reliable, research-level formalization.
@@ 1 Introduction
Recent advances in large language models (LLMs) and agentic training have revived the prospect of end-to-end formal theorem proving. On the formal side, systems such as OpenAI’s early Lean-based prover (Polu et al., 2022) and DeepMind’s AlphaProof (Hubert et al., 2025) demonstrate that reinforcement learning with formal verification feedback can produce non-trivial Lean proofs. More recently, specialized provers such as DeepSeek-Prover-V2 (Ren et al., 2025) and Seed-Prover 1.5 (Chen et al., 2025) show that large-scale agentic RL and test-time scaling can push formal success rates on established benchmarks substantially higher. These results suggest that formal proof generation is no longer limited to toy domains, and that tight tool feedback loops (retrieve–generate–verify) can be a decisive ingredient.
Figure 1: Closing the Abstraction Gap. Performance comparison of standard baselines (dashed lines) vs. our agentic approach, LeanBridge (solid lines), across difficulty levels. Bold lines highlight the best-performing models (Claude-Opus-4.5 for pass@4 baseline, GPT-5.2 for LeanBridge with NL+Statement input), while faint lines represent other models. Note the sharp performance collapse of baselines on Medium/High tasks (the “abstraction gap”), which LeanBridge effectively bridges, achieving the first successful proofs on High-difficulty problems.
@@
Despite steady progress in neural theorem proving, current formal benchmarks still do not adequately exercise library-grounded, abstraction-heavy reasoning. Widely used datasets such as miniF2F (Zheng et al., 2022) and FIMO Liu et al. (2023) largely draw from olympiad-style problems, while university-level suites like ProofNet Azerbayev et al. (2022) and PutnamBench Tsoukalas et al. (2024) focus on undergraduate competition or textbook material. These benchmarks are valuable, but they often reward short clever tricks or computation rather than sustained work inside rich abstract frameworks. By contrast, modern research mathematics is written at high generality, organized around reusable interfaces, and deeply intertwined with large libraries of definitions and lemmas—so success hinges less on a single key insight and more on navigating abstraction, managing definitions, and coherently composing library knowledge over long horizons.
@@
Category theory provides a natural stress-test for this capability: as the interface language of modern mathematics—categories, functors, natural transformations, adjunctions, limits/colimits, monads—formal proofs typically rely on diagrammatic and universal-property reasoning, i.e., constructing morphisms with the right naturality/uniqueness guarantees and showing commutativity across families of structures. Yet existing formal benchmarks rarely target this abstraction layer directly. To bridge this gap, we introduce LeanCat [^1] The benchmark code and data are available at https://github.com/sciencraft/LeanCat., a benchmark of 100 category-theory problems formalized in Lean 4 (with Mathlib), designed to test whether automated provers can operate inside a mature library and compose high-level abstractions rather than solve isolated puzzles. LeanCat complements algebra-focused benchmarks such as the FATE series Jiang et al. (2025) by shifting the frontier from abstract algebra to category theory.
Our baseline evaluation reveals a stark abstraction gap: across five strong models, the best achieves only $8.25\%$ at pass@1 and $12\%$ at pass@4. More critically, we observe a collapse in compositional generalization: performance drops from $55.0\%$ on Easy tasks to 0.0% on High-difficulty tasks (Figure 1), validating that current LLMs struggle to manage long-horizon abstraction dependencies.
To address this challenge, we go beyond static evaluation and introduce LeanBridge, a retrieval-augmented agent designed to close the abstraction gap. LeanBridge implements a dynamic retrieve–generate–verify loop that actively queries the library for definitions and lemmas. Our experiments show that LeanBridge achieves a success rate of $24.0\%$. As shown in Figure 1, our agentic approach (solid lines) effectively mitigates the performance collapse observed in baselines, securing the first successful proofs on High-difficulty tasks. These results empirically demonstrate that iterative refinement and library retrieval are not merely optimizations, but strict necessities for neuro-symbolic reasoning in abstract domains.
@@
To our knowledge, LeanCat is the first installment of a benchmark series dedicated to category theory. Beyond merely ranking models, we view this direction as pivotal for both human mathematics—by mapping which parts of abstract reasoning remain elusive to formalization—and AI—by forcing progress on abstraction-aware planning, definition retrieval, and robust compilation-driven refinement.
@@
  • LeanCat benchmark (1-Categories). We present LeanCat, a suite of 100 category-theory problems formalized in Lean 4. The tasks span eight topic clusters, curated to cover reusable abstract interfaces and structural reasoning rather than contest-style tricks.
  • Difficulty annotation pipeline. We propose a hybrid grading process combining model-based estimates and expert judgment. Each problem receives weighted ratings from both advanced LLMs and human formalizers to assign reliable Easy/Medium/High labels.
  • Baseline evaluation. We benchmark state-of-the-art generalist models (e.g., GPT-5.2, Claude-Opus-4.5) alongside specialized neural provers (e.g., DeepSeek-Prover-V2). Our extensive evaluation reveals a severe abstraction gap: while generalist models struggle with long-horizon coherence, specialized provers fail to generalize beyond their training curriculum, achieving near-zero success on Medium/High tasks even with high sampling budgets.
  • Search-augmented proving via LeanBridge. To address the abstraction gap, we propose and evaluate LeanBridge, an agentic workflow integrating LeanExplore (Asher, 2025). We demonstrate that this retrieval-augmented approach doubles the pass rate of static baselines, validating the necessity of dynamic tool use in higher mathematics.
@@ 2.1 Benchmark Structure and Content
  • Abstract part: Problems are primarily drawn from standard, widely used textbooks in category theory, especially Category Theory in Context (Riehl, 2017) and Categories for the Working Mathematician (Mac Lane, 1998), together with a small number of problems adapted from unpublished lecture notes (Kong, ; Zheng, ).
  • Concrete part: Problems are curated mainly from Abstract and Concrete Categories (Adámek et al., 1990), which provides a systematic source of exercises on concretizability, injectivity, and related themes.
  • Others: Beyond these core sources, we also include selected problems inspired by research papers and advanced community-driven references (Chen, 2021; Adámek et al., 2021).
Each LeanCat problem is self-contained at the statement level: the formal statement of the theorem is provided (often with informal description provided in the corresponding Markdown file), and all necessary definitions exist in Lean’s environment (either already in Mathlib (mathlib Community, 2020) or introduced as part of the problem setup). Where possible, we drew on known theorems from category theory literature, many tasks were crafted or adapted specifically to probe corner cases and interface interactions that an AI prover might struggle with. In several High-difficulty cases, relevant lemmas were not readily available, forcing human formalizers to devise intermediate results. This aspect of LeanCat, that it sometimes goes beyond the library, makes it a particularly stringent test for automated provers, which cannot rely solely on rote application of existing library facts.
LeanCat comprises 100 theorem statements in category theory, each fully formalized in Lean 4 (i.e. each problem is given as a Lean theorem declaration, with required definitions and context available). The problems are organized into eight topical clusters reflecting key areas of category theory, ranging from Basic Category Properties to advanced structures like Abelian Categories and Monads.
Figure 2: Sunburst diagram showing the distribution of our LeanCat problem sets by topic and difficulty. The inner ring groups problems into thematic sections (Basic, Adjunction, Reflective, Concrete, Limit, Cocompletion, Abelian, Monad), while the outer ring lists individual problems, labelled by difficulty: E = Easy, M = Medium, H = High.
Figure 2 illustrates the distribution of problems across these topics and their difficulty levels. The benchmark covers both standard textbook results (e.g., adjunctions, limits) and frontier formalization tasks (e.g., cocompletions, which require new definitions). We provide a detailed breakdown of each topical cluster and the specific mathematical competencies they test in Appendix A.
@@ 2.2 Curation Workflow
LeanCat is a statement-level benchmark: each item consists of a single standalone theorem to prove (see Figure 3 in Appendix A for a concrete example), rather than a scaffolded sequence of intermediate lemmas that gradually builds toward a final goal. This design aims to evaluate general library-grounded proving capabilities—retrieval, definition management, and abstraction navigation—instead of rewarding problem-specific hint engineering.
@@
All benchmark files follow strict, uniform conventions: (i) each Lean file is strictly self-contained, relying solely on Mathlib and explicitly including any definitions or lemmas required by the theorem statement; (ii) the natural-language problem description is stored in a companion Markdown file (supporting ), designed to be injected into the Lean code as a comment during the evaluation phase.
@@ 2.3 Difficulty Annotation Pipeline
  • LLM Difficulty Scoring. We employ a data-driven approach where each problem is evaluated using the LeanBridge sequential refinement protocol (as detailed in Section 3.1) across all five baseline models. We assign 2 points for a full proof and 0.5 points for a correct statement (credited only if the proof fails). The difficulty is calculated as: $\text{Diff}=\max(0,10-2N_{\text{proof}}-0.5N_{\text{stmt}})$, where $N_{\text{proof}}$ is the number of models producing a verified proof, and $N_{\text{stmt}}$ is the number of models generating a correct formal statement but failing the proof (ensuring disjoint counts). This scales from 0 (all solved) to 10 (none solved).
  • Human Difficulty Scoring. In parallel, two human mathematicians (with Lean expertise and category theory background) independently rated each problem on the same 1-10 difficulty scale. They factored in things like the length of the proof, intricacy of arguments, and whether any non-obvious lemmas are required. The human scores tended to correlate with intuition: e.g. a trivial diagram chase might be 2/10, whereas a complex construction spanning several definitions could be 9/10.
  • Aggregation. We combined the scores giving 50% weight to human ratings and 50% to LLM ratings. Finally, we mapped the numeric scores to difficulty categories by selecting thresholds specifically to enforce a distribution of 20 Easy, 40 Medium, and 40 High tasks. This resulted in cut-offs of scores $\leq 5.54$ for Easy and $\geq 7.8$ for High. We intentionally adopted this skewed distribution (heavier on Medium/High tasks) to prevent premature saturation by current models—which already solve $50\%$ of Easy tasks—and to provide sufficient headroom for tracking future progress in closing the abstraction gap. The complete scoring data for all 100 problems is provided in Appendix Figure 5.
This data-driven approach provides a more objective difficulty metric than human intuition alone. Our experimental results (Section 4) validate this ranking: the best-performing model solved most tasks classified as Easy, while problems with scores $\geq 9$ remained unsolved across all baselines. This distribution confirms that our difficulty labels effectively proxy the "formalization hardness" for current neuro-symbolic systems.